HIGH RESOLUTION INFRARED SPECTRA OF ^{120}SnD_{4}

\underline{M. HALONEN}^{a}, L. HALONEN^{a}, Department of Chemistry, Princeton University, NJ 08544; H. BÜRGER AND W. JERZEMBECK, FB9- Anorganische Chemie, Universität-GH Wuppertal, D-42097 Wuppertal, Germany.

High resolution FTIR spectra of the ν_{2}/ν_{4} and the ν_{1}/ν_{3} bands of ^{120}SnD_{4} have been measured with a Bruker 120 HR interferometer. The sample was prepared under vacuum using a two step synthesis. The ν_{2}/ν_{4} region (425-620 cm^{−1}) was studied using a liquid He-cooled bolometer, while an MCT detector was used in the ν_{1}/ν_{3} region (1275-1455 cm^{−1}). The resolutions of the spectra are 0.0033 cm^{−1} and 0.0024 cm^{−1}, respectively. The ν_{1}/ν_{3} band system has been analyzed rotationally, and a set of molecular parameters have been obtained. The measurement of the first and second stretching overtones is in progress, too. The results will be compared with those of normal stannane, SnH_{4}. In the case of normal stannane vibrational localisation has been found to take place in stretching vibrational overtones. Thus the dynamical symmetry of stannane, a spherical top molecule, has been found to become that of a prolate symmetric top.

aPermanent address of M. and L. Halonen: Laboratory of Physical Chemistry, P.O. Box 55 (A.I. Virtasen aukio 1), FIN-00014 University of Helsinki, Finland